OpenStudy (mya123):

how to solve -6f+13=2f-11

Awolflover1 (awolflover1):

\[−6f+13−2f=2f−11−2f\]

Awolflover1 (awolflover1):

Subtract 2f from both sides

Awolflover1 (awolflover1):

\[−8f+13=−11\]

Awolflover1 (awolflover1):

now Subtract 13 from both sides −8f+13−13=−11−13

Awolflover1 (awolflover1):

\[−11−13=?\]